# A First Look at Information Highlighting in Stack Overflow Answers

### v1

@@ Abstract
Method: In this paper, we studied 31,169,429 answers of Stack Overflow. For training recommendation models, we choose CNN and BERT models for each type of formatting (i.e., Bold, Italic, Code, and Heading) using the information highlighting dataset we collected from SO answers.
Results: Our models based on CNN architecture achieve precision ranging from 0.71 to 0.82. The trained model for automatic code content highlighting achieves a recall of 0.73 and an F1 score of 0.71, outperforming the trained models for other formatting styles. The BERT models have even lower recalls and F1 scores than the CNN models. Our analysis of failure cases indicates that the majority of the failure cases are missing identification (i.e., the model misses the content that is supposed to be highlighted) due to the models tend to learn the frequently highlighted words while struggling to learn less frequent words.
@@ 1 Introduction
Technical question and answer (Q&A) sites such as Stack Overflow (SO) have become increasingly important for software developers to share knowledge and contribute to communities. Despite Stack Overflow’s success and prevalence, navigating the knowledge on it remains challenging Gottipati et al. (2011); Nadi and Treude (2020); Xu et al. (2017); Zhang et al. (2021a). The previous study shows that finding answers in long posts remains one of the challenges Nadi and Treude (2020); Xu et al. (2017). 37% of all questions on Stack Overflow have more than one answer, and the average length of an answer is 789 characters Nadi and Treude (2020). To make the posts vivid to users, Stack Overflow platform allows users to edit their posts with Markdown and HTML Overflow (2022); StackExchange (2023), so that users can leverage various formatting styles (e.g., Bold, Italic, and Code) to highlight text and direct other users’ attention toward the most important information within posts.
@@
Information highlighting has demonstrated its effectiveness in various domains Nguyen et al. (2015); Ramírez et al. (2019); Strobelt et al. (2015); Wilson et al. (2016); Wu and Yuan (2003), including in Software Engineering Sarkar (2015) (e.g., it saves the reading time of humans and helps programming novices with code comprehension with syntax highlighting). However, little is known about how information is highlighted on technical Q&A sites (e.g., Stack Overflow). For example, how prevalent is the information highlighted? What content is highlighted using different formatting styles? Understanding this could provide a landscape of the usage of information highlighting on technical Q&A sites and shed light on the information that is considered important to developers.
@@
Previous study: In our previous study Ahmed et al. (2022), we performed the first large-scale study on the five most commonly used information highlighting types, which are Bold, Italic, Code, Delete, and Heading, to understand their characteristics and what information is highlighted in the text description of SO answers with them [^1] Note that an SO answer may contain both text and code block, we focus on the information highlighting in text description.. Our analysis of 55,209,643 information highlighting instances across 14,845,929 answers on Stack Overflow reveals that: Overall, information highlighting is prevalent on SO, i.e., 47.6% (14,845,929 out of 31,169,429) of the answers use the studied formatting types to highlight information. 38.5%, 11.3%, and 7.2% of the answers use Code, Bold, and Italic, respectively, and their highlighted content is short (median length is one word). Code formatting is mainly used to highlight source code content, such as identifiers (63.5%). Code is also used to highlight content other than source code, such as Software (4.9%) and Equation (5.2%). Italic and Bold are frequently used to highlight source code content, as well as content that warns about the context where the provided solution works or does not work, updates on answers, and references to an internal or external resource.
@@
Extention: Since information highlighting is prevalent on SO, however, identifying the appropriate content to highlight can be a challenging task for SO users (typically for new users), as they must properly pinpoint the focal points and select an appropriate formatting style, which can be time-consuming and require expertise. Recommending important information to highlight in a post can help improve the presentation of the post and make the important information more visible to SO users. Therefore, in this paper, we extend our recent study Ahmed et al. (2022) by investigating the potential of recommending highlighted content automatically. To do so, we adapt two models (i.e., CNN and BERT) to automatically identify highlighted content for each type of formatting. More specially, we train NER models on each formatting (i.e., Heading, Bold, Italic, and Code) type and recommend the content to highlight. CNN models outperform Bert-based models. CNN models achieve precision values ranging from 0.71 to 0.82. The CNN models for automatic source code content highlighting achieve a recall score of 0.73 and an F1 score of 0.75, outperforming the models for other formatting types. By analyzing the failure cases, we observed that the majority of the failure cases are due to missing identification (i.e., the model misses the content that is supposed to be highlighted). Our analysis shows that the models tend to learn the frequent highlighted words while struggling to learn less frequent words (i.g., long-tail knowledge).
@@
Paper Organization. Section 2 presents the background and the related work. Section 3 explains our research questions and how we prepare our data to answer research questions. In Section 4, we present the results of our research questions. Section 5 discuss the implications of our study and threats to validity. Finally, Section 6 concludes our study.
2 Background & Related work
@@ 2.1 Background
Stack Overflow allows users to use Markdown and HTML to write and edit posts Overflow (2022); StackExchange (2023). Certain formatting types are used to highlight the information in text description with Markdown and HTML tags. In this study, we focus on the five most commonly used formatting types, which are Bold, Italic, Delete, Code, and Heading. In Table 1, we present the HTML tags and their equivalent Markdown syntax for each formatting type. We group HTML tags and Markdown formatting based on their rendering effect. For example, we group <b> and <strong> as Bold since they render the same effect in the browser when users read the posts on Stack Overflow. In this study, we consider a highlighted sequence of words as an highlighted instance. One post could have multiple highlighted instances. For instance, in Figure 1, the post has 10 Code instances [^2] https://stackoverflow.com/questions/32402475.
@@
Code
<code>
’example’
Bold
<b>, <strong>
**example**, __example__
Italic
<i>, <em>
*example*
Delete
<del>, <s>
None
Heading
<h1>, <h2>, <h3>, <h4>, <h5>, <h6>
# example, ## example, ### example, #### example, ##### example, ###### example
2.2 Related work
2.2.1 Information highlighting
Previous research has explored the benefits of highlighting information in various domains over the last decades Nguyen et al. (2015); Ramírez et al. (2019); Strobelt et al. (2015); Wilson et al. (2016); Wu and Yuan (2003). Wu and Yuan have shown that highlighting could reduce the cognitive load and thus reduce reading time Wu and Yuan (2003). Jorge et al. investigated the impact of highlighting text in the text classification tasks in the machine learning area and found that highlighting is effective in reducing classification effort for humans Ramírez et al. (2019). Similarly, Nguyen et al. developed a tool to explain the output ML models by highlighting the portion of text and demonstrated its effectiveness in model explanation Nguyen et al. (2015). Various studies have been done to investigate the impact of information highlighting in software engineering, such as code comprehension Beelders and du Plessis (2016); Hannebauer et al. (2018); Palma et al. (2022); Sarkar (2015), source code evolution Escobar et al. (2022). For instance, Advait Sarkar investigated the effect of syntax highlighting on program comprehension and its interaction with programming experience Sarkar (2015), and found that syntax highlighting significantly improves task completion time. Different from prior studies that focus on understanding the influence of information highlighting, we focus on understanding the practice of information highlighting on SO.
2.2.2 Knowledge retrieval on Stack Overflow
Several studies have investigated developers’ challenges and solutions in retrieving knowledge from technical Q&A sites and provided some tools to facilitate the process Gottipati et al. (2011); Nadi and Treude (2020); Xu et al. (2017); Zhang et al. (2021a). For instance, Gottipati et al. found that in software forums it is a painstaking process for users to manually search through answers in various threads of posts Gottipati et al. (2011), and they developed a customized search engine to find relevant answers from various software forums. Similarly, Xu et al. developed a technique called AnswerBot, which takes input as the technical question and generates an answer summary for the question Xu et al. (2017). Zhang et al. pointed out an issue of current comment ranking and display mechanism on SO (e.g., a large portion of useful comments are not visible to users) and proposed an alternative ranking mechanism to alleviate the issue Zhang et al. (2021a). Sarah and Treude investigated the possibility of using two existing techniques, wordpartten and lexrank, to identify the essential sentences from a long SO answer Nadi and Treude (2020). Different from prior studies that focus on retrieving relevant or important answers from SO, we investigate what and how information is highlighted on SO. Our study could provide insights to enhance existing techniques.
2.2.3 Recommendation system for Stack Overflow
Previous studies developed different recommendation tools to ease the use of the Q&A sites, such as recommending hyperlinks Li et al. (2016, 2018b), tags He et al. (2022); Maity et al. (2019); Rekha et al. (2014); Wang et al. (2019a, 2018), and similar questions Wang et al. (2019b); Yin et al. (2019). Li et al. developed LinkLiv to recommend hyperlinks Li et al. (2018b), which uses multiple features, including hyperlink co-occurrences in Q&A discussions, and locations (e.g., question, answer, or comment). Wang et al. developed EnTagRec, which combines two complementary lines in the statistics community Bayesian and frequentist Wang et al. (2018). Wang et al. developed SOTagRec combining a convolutional neural network and collaborative filtering model to infer tags for new postings by studying the historical postings and their tags Wang et al. (2019a). Maity et al. developed DeepTagRec Maity et al. (2019), which learns from the content representation of question title and body, and recommends appropriate question tags on Stack Overflow. Wang et al. developed an approach that combines topical interest, topical expertise, and activeness to recommend answers for new questions Wang et al. (2019b). Similar to these studies, we investigate the potential of developing models for recommending information to highlight in SO answers. We adapted Named Entity Recognition (NER) models to identify the content that needs to be highlighted.
3 Research Questions & Data preparation
@@ 3.1 Research questions
In RQ1, we aim to understand the prevalence of the usage of the studied highlighting formatting and their characteristics. Understanding this can provide practitioners an overview of the landscape of information highlighting in SO answers. We study Code formatting and the rest formatting types (i.e., Bold, Italic, Heading, and Delete) in separated RQs (RQ2 and RQ3) due to their different nature. We study the information highlighted by different formatting types. By understanding this, we can provide insights into what information is important from users’ perspectives and provide insights for future research. For instance, we can provide insight for the downstream research that leverages SO information to facilitate software engineering tasks (e.g., API documentation enrichmentLi et al. (2018a); Treude and Robillard (2016)). Identifying the appropriate content to highlight can be a challenging task for SO users, as they must properly pinpoint the focal points and select an appropriate formatting style, which can be time-consuming and require expertise. For instance, we observe that, in many SO answer posts, certain contents were not initially highlighted in the first version of the post but were highlighted in subsequent versions, such as the second or third. For instance, in the second and third revisions of a post, a user highlighted the code-related content (e.g., “Integer”, “Exception”) to improve the presentation of the post [^3] https://stackoverflow.com/posts/218510/revisions?page=2. Therefore, in RQ4, we explore the feasibility of utilizing machine learning models to recommend which information to highlight.
@@ 3.2 Data preparation
To answer the four RQs, we downloaded a data dump of Stack Overflow from the Stack Exchange data dump dated March 2021444https://archive.org/details/stackexchange. The data dump contains details information about posts (i.e., questions and answers), as well as their revision history. In this study, we include all the answers and we ended up with 52,166,061 answer posts.
@@
As the dataset was large for processing, we imported it into MySQL. For each answer post, we first extract the textual content from its body and exclude code block(s). Note that we extracted code block(s) using tag “<pre>”. Next, we apply the regular expressions defined in Table 1 to identify information highlighting in each answer post, using the Python library Regex. In the end, we extract the content that is highlighted by the studied formatting types. We make our replication package public at https://github.com/shaoweiwang2010/REP_2022_Information_highlight_SO.

### v3

# Studying and Recommending Information Highlighting in Stack Overflow Answers

@@ Abstract
Method: In this paper, we studied 31,169,429 answers of Stack Overflow. For training recommendation models, we choose CNN-based and BERT-based models for each type of formatting (i.e., Bold, Italic, Code, and Heading) using the information highlighting dataset we collected from SO answers.
Results: Our models achieve a precision ranging from 0.50 to 0.72 for different formatting types. It is easier to build a model to recommend Code than other types. Models for text formatting types (i.e., Heading, Bold, and Italic) suffer low recall. Our analysis of failure cases indicates that the majority of the failure cases are due to missing identification. One explanation is that the models are easy to learn the frequent highlighted words while struggling to learn less frequent words (i.g., long-tail knowledge).
@@ 1 Introduction
Technical question and answer (Q&A) sites such as Stack Overflow (SO) have become increasingly important for software developers to share knowledge and contribute to communities. Despite Stack Overflow’s success and prevalence, navigating the knowledge on it remains challenging [12, 26, 51, 56]. The previous study shows that finding answers in long posts remains one of the challenges [26, 51]. 37% of all questions on Stack Overflow have more than one answer, and the average length of an answer is 789 characters [26]. To make the posts vivid to users, Stack Overflow platform allows users to edit their posts with Markdown and HTML [29, 41], so that users can leverage various formatting styles (e.g., Bold, Italic, and Code) to highlight text and direct other users’ attention toward the most important information within posts.
@@
Information highlighting has demonstrated its effectiveness in various domains [27, 34, 42, 49, 50], including in Software Engineering [37] (e.g., it saves the reading time of humans and helps programming novices with code comprehension with syntax highlighting). However, little is known about how information is highlighted on technical Q&A sites (e.g., Stack Overflow). For example, how prevalent is the information highlighted? What content is highlighted using different formatting styles? Understanding this could provide a landscape of the usage of information highlighting on technical Q&A sites and shed light on the information that is considered important to developers.
@@
Previous study: In our previous study [2], we performed the first large-scale study on the five most commonly used information highlighting types, which are Bold, Italic, Code, Delete, and Heading, to understand their characteristics and what information is highlighted in the text description of SO answers with them [^1] Note that an SO answer may contain both text and code block, we focus on the information highlighting in text description.. Our analysis of 55,209,643 information highlighting instances across 14,845,929 answers on Stack Overflow reveals that: Overall, information highlighting is prevalent on SO, i.e., 47.6% (14,845,929 out of 31,169,429) of the answers use the studied formatting types to highlight information. 38.5%, 11.3%, and 7.2% of the answers use Code, Bold, and Italic, respectively, and their highlighted content is short (median length is one word). Code formatting is mainly used to highlight source code content, such as identifiers (63.5%). Code is also used to highlight content other than source code, such as Software (4.9%) and Equation (5.2%). Italic and Bold are frequently used to highlight source code content, as well as content that warns about the context where the provided solution works or does not work, updates on answers, and references to an internal or external resource.
@@
Extention: Since information highlighting is prevalent on SO, however, identifying the appropriate content to highlight can be a challenging task for SO users (typically for new users), as they must properly pinpoint the focal points and select an appropriate formatting style, which can be time-consuming and require expertise. Recommending important information to highlight in a post can help improve the presentation of the post and make the important information more visible to SO users. Therefore, in this paper, we extend our recent study [2] by investigating the potential of recommending highlighted content automatically. To do so, we adapt two models (i.e., CNN and BERT) to automatically identify highlighted content for each type of formatting. More specially, we train NER models on each formatting (i.e., Heading, Bold, Italic, and Code) type and recommend the content to highlight. CNN-based models outperform BERT-based models. CNN models achieve precision values ranging from 0.50 to 0.72. The CNN models for automatic source code content highlighting achieve a recall score of 0.66 and an F1 score of 0.69, outperforming the models for other formatting types. In addition, CNN-based models outperform BERT-based models. By analyzing the failure cases, we observed that the majority of the failure cases are due to missing identification (i.e., the model misses the content that is supposed to be highlighted). Our analysis shows that the models tend to learn the frequent highlighted words while struggling to learn less frequent words (i.g., long-tail knowledge). We make our replication package public at https://github.com/shaoweiwang2010/REP_2022_Information_highlight_SO.
@@
Paper Organization. Section 2 presents the background and motivation of the work. Section 3 explains our research questions and how we prepare our data to answer research questions. Section 4 introduces the methodology of our study. In Section 5, we present the results of our research questions. Section 6 discusses the implications of our study and threats to validity. Section 7 presents the related work. Finally, Section 8 concludes our study.
2 Background
@@ 2.1 Background
Stack Overflow allows users to use Markdown and HTML to write and edit posts [29, 41]. Certain formatting types are used to highlight the information in text description with Markdown and HTML tags. In this study, we focus on the five most commonly used formatting types, which are Bold, Italic, Delete, Code, and Heading. In Table 1, we present the HTML tags and their equivalent Markdown syntax for each formatting type. We group HTML tags and Markdown formatting based on their rendering effect. For example, we group <b> and <strong> as Bold since they render the same effect in the browser when users read the posts on Stack Overflow. In this study, we consider a highlighted sequence of words as an highlighted instance. One post could have multiple highlighted instances. For instance, in Figure 1, the post has 10 Code instances [^2] https://stackoverflow.com/questions/32402475.
@@
Code <code> ’example’
Bold <b>, <strong> **example**, __example__
Italic <i>, <em> *example*
Delete <del>, <s> None
Heading <h1>, <h2>, <h3>, <h4>, <h5>, <h6> # example, ## example, ### example, #### example, ##### example, ###### example
2.2 Motivation
As discussed in Section 2, Stack Overflow allows users to apply different types of formatting on content to highlight the information. By comprehending the usage practices of those formatting types, we can provide insights into what information is important from users’ perspectives and provide insights for future research. For instance, we can provide insight into the downstream research that leverages SO information to facilitate software engineering tasks (e.g., API documentation enrichment[20, 44]). Furthermore, identifying the appropriate content to highlight can be a challenging task for SO users, as they must properly pinpoint the focal points and select an appropriate formatting style, which can be time-consuming and require expertise. For instance, we observe that, in many SO answer posts, certain contents were not initially highlighted in the first version of the post but were highlighted in subsequent versions, such as the second or third. For instance, as an example shown in Figure 2, in the second and third revisions of a post, a user highlighted the code-related content (e.g., “Integer”, “Exception”) to improve the presentation of the post. In addition, we observed (see details in Sections 5.2 and 5.3). Therefore, in this study, we also aim to investigate the feasibility of building automatic approaches to recommend content to be highlighted with formatting types.
Figure 2: An example of revising highlight to improve the presentation of the post.
3 Experimental settings
@@ 3.1 Research questions
In RQ1, we aim to understand the prevalence of the usage of the studied highlighting formatting and their characteristics. We study Code formatting and the rest formatting types (i.e., Bold, Italic, Heading, and Delete) in separated RQs (RQ2 and RQ3) due to their different nature. We study the information highlighted by different formatting types. In RQ4, we explore the feasibility of utilizing machine learning models to recommend which information to highlight.
@@ 3.2 Data preparation
To answer the four RQs, we downloaded a data dump of Stack Overflow from the Stack Exchange data dump dated March 2021333https://archive.org/details/stackexchange. The data dump contains details information about posts (i.e., questions and answers), as well as their revision history. In this study, we include all the answers and we ended up with 31,169,429 answer posts.
@@
As the dataset was large for processing, we imported it into MySQL. For each answer post, we first extract the textual content from its body and exclude code block(s). Note that we extracted code block(s) using tag “<pre>”. Next, we apply the regular expressions defined in Table 1 to identify information highlighting in each answer post, using the Python library Regex. In the end, we extract the content that is highlighted by the studied formatting types.
Table 2 presents the basic statistics of our dataset for different formatting types.
4 Methodology
Figure 3 presents an overview of the methodology for our study. Particularly, our study could be decomposed into two parts. In part 1, we perform an empirical study on the information highlighting of SO posts. In part 2, we investigate the feasibility of building a recommendation system for information highlighting. We elaborate on the details of the approach to answering RQs below.
Figure 3: Framework of our study.
4.1 Approach for RQ1
To understand the prevalence of information highlighting in SO answers, we compute the percentage of answers that have the studied formatting types (see Table 1) and the percentage of words highlighted in each answer with each formatting. In addition, we compute the basic statistics of information highlighting instances to understand the characteristics of each formatting type. More specifically, we calculate the distribution of instances for each formatting and the number of words highlighted with them.
4.2 Approach for RQ2
In this RQ, we aim to understand what content is highlighted with the Code. First, we randomly sampled 385 Code instances with a 5% interval and 95% confidence level. Since there is no existing terminologies we could reuse for this purpose, we manually performed a lightweight open coding-like process [38, 55] to identify the type of content highlighted with Code. The process involves three phases and is performed by the first authors (A1 and A2) of the paper. In phase I, A1 and A2 derived a draft list of types based on 50 Code instances. During the phase, the types were revised and refined. In phase II, A1 and A2 independently applied the derived types to the rest 335 samples. They took notes regarding the diffidence or ambiguity during the labeling. During this phase, no new types were introduced. In phase III, A1 and A2 discussed the results from Phase II to resolve any disagreements until a consensus was reached. The coding process has a Cohen’s kappa of 0.8 (measured before starting Phase III), which indicates a substantial level of the inter-rater agreement [45]. Table 5 presents the definition and the corresponding example of the derived types for Code.
4.3 Approach for RQ3
In this RQ, we aim to understand what content is highlighted with the studied Text formatting (i.e., Bold, Italic, Delete, and Heading). Similar to RQ2, we performed an manual study. We randomly sampled 385 Text instances with a 5% interval and 95% confidence level. We ended up with 177 Bold, 169 Italic, 3 Delete, and 36 Heading instances. We then identified the types of highlighted content by following the methodology used in RQ2. The coding process has a Cohen’s kappa of 0.78 (measured before starting Phase III), which indicates a substantial level of the inter-rater agreement [45]. Table 6 presents the definition and the corresponding example of the derived types for Text formatting.
4.4 Approach for RQ4
In the previous RQs, we focused on analyzing the highlighted content on Stack Overflow and we noticed that users misuse formatting types. For instance, users used Bold and Italic to highlight code. Therefore, we aim to explore the feasibility of implementing an automated machine learning-based method for content highlighting. This feature would be helpful, particularly for SO users who are inexperienced in the practice of highlighting information and seek to automatically highlight critical content when composing their posts. Specifically, we aim to create a model that can highlight information in an input answer post using appropriate formatting types (Bold, Italic, Delete, Code, and Heading). To achieve this, we use neural network architectures initially designed for Named Entity Recognition (NER) without actually training NER models. Our task shares similarities with NER, such as the short length of labeled (highlighted) tokens and a limited number of text labeling (formatting) options. Moreover, both NER and automated information highlighting can be formed as a sequence labeling task. Note that we train the neural networks from scratch on our information highlighting dataset.
Figure 4: Layers of the CNN model
In this RQ, we have used two different structured models to investigate. One model is based on the Convolutional Neural Network (CNN) model and another is based on the transformer model, BERT [8]. We selected CNN because it has been widely used in NER tasks to capture the context information for predicting NER labels and achieved good performance [7, 58, 16]. Compared to Recurrent Neural Network-based approaches (RNN), CNN-based approaches are more efficient. In recent years, the emergence of pre-trained models, such as BERT [8], T5 [32], GPT [5] have re-evolute the landscape of the natural language processing (NLP) domain and demonstrated their effectiveness in NLP tasks, including NER tasks. Pre-trained models encapsulate knowledge gleaned from a massive amount of training data, bypassing the need for training from scratch, and could be easily adapted to downstream tasks by fine-tuning. BERT is an encoder-based model and considers the contexts of both preceding and subsequent words. BERT has been proven effective in embedding the context for NER task [39, 6, 13]. Therefore, we also investigate BERT in this RQ.
Figure 4 presents the architecture of our CNN model. In this paper, we use the lib spaCy [40] as the implementation of CNN model. We opt to utilize spaCy since it has emerged as the de facto standard for practical NLP due to its speed, robustness, and performance [19]. The model architecture comprises a token embedding layer (128 dimensions), four convolutional layers, and an output layer. The token embedding layer is named as MultiHashEmbed. It features an embedding layer that embeds a number of lexical attributes separately. Then the results are concatenated as a vector semantic representation of a token for passing through a feed-forward network. As it uses hashing to store embeddings in memory, it also supports faster processing of large datasets. The embedding layer is followed by a 4-layered Convolutional Neural Network (CNN), named MaxoutWindowEncoder. Each of the convolutional layers is followed by a max pooling layer. The final layer is a softmax layer for predicting labels on tokens. Training employs cross-entropy loss and the Adam optimizer.
Figure 5: Layers of the NER BERT model
Figure 5 presents the architecture of BERT model, which consists of multiple transformer blocks, each containing a self-attention mechanism for getting the context of the input sequence. In this work, we fine-tuned BERT for our prediction task. We used the most recent checkpoint of BERT model from HuggingFace platform [11].
Dataset preparation: Prior to constructing the training and testing datasets, we carefully cleaned the dataset to eliminate misuse instances (e.g., using bold formatting to highlight code) to ensure that the dataset used for model training would accurately represent the intended formatting of information. To achieve this, we implemented the following pre-processing steps to clean the data.
  • 1.
    Cleaning up code dataset. According to Stack Overflow guidelines, users are encouraged to highlight code content using the code tag exclusively. However, our analysis of RQ3 results revealed instances where code tag was improperly used for Path, software, Terminology, and Equation. Therefore, our first task is to remove such cases from our code dataset. We used regular expressions (e.g., ^(/[^/ ]*)+/?\$) to identify all paths and equations in content highlighted with code. For identifying software and terminologies, we collected all SO tags (e.g., “mysql”, “andriod”, “django”, and “hashtable”) [^4] https://stackoverflow.com/tags as our comprehensive dictionary. We then used fuzzy matching to identify such software and terminologies in the code dataset based on the dictionary. We used fuzzywuzzy library [^5] https://pypi.org/project/fuzzywuzzy/ for our implementation. The reason we selected SO tags as our dictionary is that they contain a diverse and wide range of software names and technologies [53]. we discarded those cases from our code dataset.
  • 2.
    Cleaning up text dataset. Users sometimes highlight code data with text formatting tags. As it goes against the SO guidelines, we need to identify the code that is highlighted with text formatting tags (i.e., Bold, Italic, Heading). To identify the code content in text dataset, we cross-checked the code dataset and text dataset and examined if any content was cross-highlighted in both code and text formatting. We removed such instances from text data. For instance, if we find “BufferedReader.close()” are both highlighted with Code and Bold formatting, we only keep the instances in code data and remove the instances from text data.
  • 3.
    Removing other HTML tags. We used regular expressions to remove other HTML tags rather than the targeting tags. For example, for the Bold model, we only kept the <b> and <strong> tags to identify the bold highlighting in sentences and removed other highlighting tags.
Now we can construct our training and testing set. We divided the post into sentences and extracted the corresponding formatting tags described in Section 3. In other words, each data point is a highlighted instance (a sentence containing at least one type of information highlighting). When preparing the training data, we consider the content highlighted with different formatting types as the entities in the NER model. We annotate the tagged words with their formatting tags, considering their position in the phrase. For example, in the sample highlighted sentence shown in Figure 4, i.e., “You will have to use client_wait_for().”, “client_wait_for()” is formatted with a code tag. To label this sentence, we assign the label “B-code” to the highlighted word, indicating that it is the beginning of a highlighted Code content, while the other words in the sentence are labeled as O (Outside). If multiple words are formatted, we used I (Interior) and E (Ending) to label the highlighted content’s internal words and the ending word. For example, “you need to run.git push --force --all first” contains more than one word in the highlighted Code content. So the labels for the phrase would be (B-code: “git”), (I-code: “push”), (I-code: “--force”) and (E-code: “--all”). Note that we train separate models for each highlighting type. When preparing the dataset for a particular highlighting type, if a sentence has other highlighting types rather than the one we target, we omit other types of highlighting types from the sentence.
Table 2: Number of sentences in train and test datasets.
Type Training set Testing set
Bold 521308 140539
Italic 477690 104128
Code 1410890 406960
Heading 131976 21028
Following prior studies, we partitioned the dataset into an 80:20 ratio for training and testing purposes [31, 3, 52]. We trained our models in 3 epochs where the learning rate is 0.001. We employ two training settings.
We use 32 as the size of the batch in the fixed-size setting (i.e., fixed). Models based on BERT architecture are trained in 0.05 learning rate using 16 batches(fixed) in 3 epochs.
Table 3: Models parameters.
Parameter CNN-based BERT-based
Epochs 3 3
Learning rate 0.001 0.05
Batch size 32 16
Evaluation: We use precision, recall, and F1-score to measure the performance of models, following prior studies evaluating NER models [10, 17]. Instead of using the exact match (i.e., a true positive is counted when all of the tokens of a named entity are labeled correctly, both the length and type), we decide to use a partial match [25, 43], in which we consider if each individual token is predicted correctly. The partial match allows us to measure the performance of the segmentation aspects of highlighted content when the content has multiple words. More specifically, we calculate the precision and recall as follows: